The average annual wind speed in Rochester, Minnesota is 15.5 miles per hour. If a sample of 95 days are used to determine the average wind speed, find the 98% confidence interval of the mean. Assume the standard deviation was 3.2 miles per hour. (Show Work)
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 15.5

Population standard deviation =    = 3.2

Sample size = n = 95

At 98% confidence level

= 1 - 98%  

= 1 - 0.98 =0.02

/2 = 0.01

Z/2 = Z0.01 = 2.33


Margin of error = E = Z/2 * ( /n)

= 2.33 * ( 3.2 /  95 )

= 0.76

At 98% confidence interval estimate of the population mean is,

  ± E

15.5 ± 0.76

( 14.74, 16.26 )